OpenStudy (anonymous):

Determine if the conditional is true or false: If two numbers are odd, then their sum is even.

OpenStudy (anonymous):

true. ex. 2+3=5 1+3=4

OpenStudy (anonymous):

A number is odd if x = 2n + 1 for any n in Z (set of integers). A number is even if x = 2n for any n in Z. so if a,b are odd. then a = 2n1 + 1 ==> n1 is n subscript 1 b = 2n2+ 1 ==> n2 is n subscript 2 then a+b = 2n1 + 2n2 +2 = 2(n1+n2+1) . n1 + n2 + 1 is inside Z since n1 and n2 is inside Z hence a+b is written as 2k which is even. k= n1 + n2 +1

OpenStudy (anonymous):

TRUe